# Robin3D[Uncaptioned image]  : Improving 3D Large Language Model via Robust Instruction Tuning

### v1

@@ Abstract
Recent advancements in 3D Large Language Models (3DLLMs) have highlighted their potential in building general-purpose agents in the 3D real world, yet challenges remain due to the lack of high-quality robust instruction-following data, leading to limited discriminative power and generalization of 3DLLMs. In this paper, we introduce Robin3D, a powerful 3DLLM trained on large-scale instruction-following data generated by our novel data engine, Robust Instruction Generation (RIG) engine. RIG generates two key instruction data: 1) the Adversarial Instruction-following data, which features mixed negative and positive samples to enhance the model’s discriminative understanding. 2) the Diverse Instruction-following data, which contains various instruction styles to enhance model’s generalization. As a result, we construct 1 million instruction-following data, consisting of 344K Adversarial samples, 508K Diverse samples, and 165K benchmark training set samples. To better handle these complex instructions, Robin3D first incorporates Relation-Augmented Projector to enhance spatial understanding, and then strengthens the object referring and grounding ability through ID-Feature Bonding. Robin3D consistently outperforms previous methods across five widely-used 3D multimodal learning benchmarks, without the need for task-specific fine-tuning. Notably, we achieve a 7.8% improvement in the grounding task (Multi3DRefer) and a 6.9% improvement in the captioning task (Scan2Cap). Code is available at https://github.com/WeitaiKang/Robin3D.
@@ 1 Introduction
Spatial Intelligence [Li, 2024] refers to the ability of AI to understand the 3D world and reason within 3D space. Related ideas, such as Embodied AI [Duan et al., 2022] and Robotic Agent [Bousmalis et al., 2023], express a similar aim to build general-purpose assistants in the 3D real world. To achieve this goal, researchers have drawn inspiration from the success of 2D Multimodal Large Language Models (MLLMs) [Liu et al., 2024c; You et al., 2023] and have started exploring the potential of 3D Large Language Models (3DLLMs) to create general agents [Hong et al., 2023; Chen et al., 2024a; Wang et al., 2023b; Huang et al., 2023a] in the 3D domain or to attain Spatial Intelligence.
Instruction-following tuning [Liu et al., 2024c; a; b] in MLLMs refers to training the LLM to execute natural language commands by integrating both textual and visual information. In contrast to the versatile image-text pairs employed for training 2D MLLMs, collecting 3D instruction-following data for 3DLLM remains a significant challenge. Although existing works have made progress [Hong et al., 2023; Huang et al., 2023b; Chen et al., 2024c] in generating more instruction data, they still lack robustness in two aspects: 1) Most of the existing instruction data consist of positive pairs, lacking adversarial or negative samples. Therefore, models trained on such data tend to be less discriminative because they might overfit to the positive pairs and are more likely to hallucinate positive responses to any input. 2) Some instruction data also lack diversity in language styles, as human annotators or generative models [OpenAI, ; Wang et al., 2023a] are typically asked to follow fixed instructions when describing objects [Chen et al., 2020; 2021], or the data is generated using predefined templates [Achlioptas et al., 2020], which may limit models’ generalizability.
To address these challenges, we introduce Robin3D, a robust and powerful 3D Large Language Model tuned on large scale instruction-following data generated by our novel data engine, Robust Instruction Generation (RIG) engine. Specifically, RIG is designed to generate two types of data: Adversarial Instruction-following data, which is characterized by mixing adversarial or negative samples with positive ones. This process decouples the potential memorized positive pairs in the training set, leading to a more discriminative understanding of individual objects and instructions. To present a comprehensive adversarial dataset, we cover both object-level and scene-level instructions, from category-based identification problems to expression-based reasoning challenges, resulting in four new tasks. Diverse Instruction-following data, which first comprehensively collects various types of instructions from existing studies or transforms current tasks into instruction-following format. To harness the powerful in-context learning capability of LLMs, we use ChatGPT [OpenAI, ] to diversify the language styles of the instructions by crafting specific prompt engineering templates tailored to each task. Combining these with the original training sets of current benchmarks, we construct 1 million instruction-following samples, with approximately 344K adversarial data (34%), 508K diverse data (50%), and 165K benchmark data (16%), as shown in Fig. 1 (right).
@@
Figure 1: Robin3D surpasses previous SOTA across all the benchmarks (left), by training on our RIG-generated 1 million data (right).
Akin to current SOTA approaches [Huang et al., 2023a; ], our Robin3D model uses pre-trained 3D [Zhou et al., 2023] and 2D [Oquab et al., 2023] models to effectively represent 3D scenes at the object level, leveraging object IDs [Huang et al., 2023a] to refer to the objects in the input or ground them in the output. However, the original methods inevitably weaken the spatial relationships between object features due to the object-centric normalization in their 3D backbone [Zhou et al., 2023], which hinders the learning of our diverse visual grounding data. Moreover, in the current approaches [Huang et al., 2023a; ], object IDs are simply appended to object features in an interleaved manner, which may lack sufficient connection between the IDs and the features, making it struggle with complex referring and grounding requirements in our adversarial instruction data. To tackle these problems, Robin3D first integrates the Relation-Augmented Projector (RAP) to enhance spatial relationship understanding by enriching object-centric features with scene-level context and positional information. We then introduce ID-Feature Bonding (IFB), which bonds each ID to its corresponding features by wrapping the features with identical ID tokens. IFB further reinforces this connection by a post-vision order, resulting in a more informative ID-feature sequence to improve the referring and grounding ability.
We conduct comprehensive experiments across five representative 3D multimodal learning benchmarks, including ScanRefer [Chen et al., 2020], Multi3DRefer [Zhang et al., 2023], Scan2Cap [Chen et al., 2021], ScanQA [Azuma et al., 2022], and SQA3D [Ma et al., 2022]. As shown in Fig. 1 (left), our Robin3D achieves state-of-the-art performance across all the benchmarks without fine-tuning on specific tasks, making a significant step towards Spatial Intelligence.
In summary, our contributions are threefold: (i) We introduce Robin3D, a powerful 3DLLM trained on robust instruction-following data generated by our novel data engine, RIG. (ii) We incorporate two modules into Robin3D, RAP and IFB, to enhance its spatial understanding as well as its referring and grounding capabilities when handling our complex instructions. (iii) Our Robin3D achieves state-of-the-art performance across five widely-used benchmarks without task-specific fine-tuning.
@@ 2 Related Work
Recent advancements in 3D vision-language (3D-VL) learning [Chen et al., 2020; Achlioptas et al., 2020; Azuma et al., 2022; Chen et al., 2021; Kang et al., 2024b] have focused on bridging the gap between 3D scene understanding and natural language, which is essential for developing embodied intelligence. Similar to 2D vision-language learning [Kazemzadeh et al., 2014; Kang et al., 2024c; Johnson et al., 2016; You et al., 2023; Kang et al., 2024a; Antol et al., 2015; Zhang et al., 2024; Kang et al., 2024d], a variety of tasks such as 3D Visual Grounding [Chen et al., 2020; Zhang et al., 2023; Achlioptas et al., 2020], 3D Dense Captioning [Chen et al., 2021], and 3D Question Answering [Azuma et al., 2022; Ma et al., 2022] have been proposed to evaluate the ability of human instruction following in relation to 3D object properties and spatial relationships. Initial efforts focus on building a task-specific model for a single tasks, such as EDA [Wu et al., 2023] for grounding and Vote2Cap-DETR++ [Chen et al., 2024b] for captioning. Recent research has shifted towards developing unified models capable of handling multiple 3D scene-language tasks. Approaches like 3DJCG [Cai et al., 2022] and D3Net [Takahashi & Mitsufuji, 2020] leverage task synergies, while 3D-VisTA [Zhu et al., 2023], 3D-VLP [Jin et al., 2023] and PQ3D [Zhu et al., 2024] introduce pre-training techniques and unified representations to align 3D visual data with language. However, their dependence on task-specific heads restricts their flexibility for more generalized user-assistant interactions.
@@
Following the success of 2D Multimodal Large Language Models [You et al., 2023; Zhang et al., 2024; Liu et al., 2024c; a; b], researchers begin to explore the reasoning abilities and promising generalizability of LLMs in human instruction following within the 3D domain, leading to the emergence of 3D Large Language Models (3DLLMs). Models like PointLLM [Xu et al., 2023] and Imagebind-LLM [Han et al., 2023] show strong performance in object-level tasks by mapping 3D data into LLMs. However, they face difficulties in handling scene-level reasoning. 3D-LLM [Hong et al., 2023] incorporates positional embeddings and location tokens, and Oryx [Liu et al., 2024d] offers a solution to support multi-view arbitrary resolution image. However, their reliance on 2D encoders limits its ability to fully capture 3D spatial structures. Models such as LL3DA [Chen et al., 2024a], Chat-3D [Wang et al., 2023b], LEO [Huang et al., 2023b], and Scene-LLM [Fu et al., 2024] have made progress in improving scene-level dialogue capabilities, showing promising results in question-answering and captioning tasks. However, their insufficient visual grounding capability limits their application in Embodied AI or Robotic Agents, which require precise object localization and manipulation following human instruction. To further enhance grounding abilities, Grounded 3D-LLM [Chen et al., 2024c] introduces referent tokens and contrastive learning to unify grounding and textual responses. Similarly, Chat-3D v2 [Huang et al., 2023a] proposes the use of object identifiers (object IDs) for referring and grounding. Building on Chat-3D v2[Huang et al., 2023a], Chat-Scene [Huang et al., ] further incorporates DINO v2 [Oquab et al., 2023] to provide strong multi-view, object-centric 2D representations. Despite these advancements, current 3D LLMs, which are trained solely on positive 3D visual-language pairs and template-based instructions, suffer from suboptimal generalization and a potential for overfitting.
3 Methodology
@@ 3.1 Preliminary
To train a 3D LLM using instruction fine-tuning, we first represent the 3D scene as a sequence of vision tokens, then append it with system and question prompts, expressed as sequences of language tokens, to indicate the task. Taking the above tokens as input, a LLM is supervised to output the answer tokens via next token prediction. Specifically, as shown in Fig. 2, given the point cloud of a 3D scene, we use the pre-trained 3D segmenter Mask3D [Schult et al., 2023] to extract object features along with their corresponding 3D masks. Following Huang et al. , we further sample each object’s point cloud based on the 3D masks, normalize it, and employ the pre-trained Uni3D [Zhou et al., 2023] to extract unified object-centric 3D features. Additionally, 2D masks projected from the 3D masks are used to sample and average 2D features, which are extracted by DINO v2 from multi-view images of each object. Our Relation-Augmented Projector fuses the 3D features and position embeddings from Mask3D and Uni3D into our final 3D features. In line with Huang et al. , we incorporate special tokens $\{<\texttt{OBJ}_{i}>\}_{i=1...n}$ as object IDs into the vocabulary. These ID tokens are paired with 2D & 3D object features to indicate each object, for referring to the object in the input instruction or grounding the object in model’s output. Our ID-Feature Bonding combines each object feature with its corresponding object ID, and appends the system and question prompts at the beginning of the sequence, which are then fed into the LLM. For more details on the object IDs and the extraction of different types of features, please refer to Huang et al. .
Figure 2: Overview of the proposed Robin3D model structure. Bottom: Our Relation-Augmented Projecter fuses the features and position embedding from Mask3D and Uni3D to generate final 3D features. 2D features from DINO v2 are projected into the LLM space. We freeze the Mask3D, Uni3D, and DINO v2. Middle: ID-Feature Bonding enhances the connection between object IDs and object features by wrapping the features with identical IDs and the Post-Vision order. Top: We use LoRA to fine-tune the LLM on our constructed 1 million instruction data.
3.2 Relation-Augmented Projecter (RAP)
As shown in Fig. 2 (bottom), to obtain relation-aware 3D features while preserving the unified object-centric characteristics, our RAP considers three types of 3D features: a) the object features from Mask3D, $X_{mask3d}$, which are scene-level respresentations, containing spatial relationship, as they come across multiple cross-attention layers to exchange information. b) the position embeddings of the Mask3D, $X_{pos}$, which are directly projected from the coordinates of the corresponding object features. c) the unified object features, $X_{uni3d}$, from Uni3D. Our RAP is formulated as:
$X=\text{Concat}(\text{Norm}_{L2}(X_{uni3d}),\text{Norm}_{L2}(X_{mask3d})),\quad X_{rap}=\text{MLP}(X)+\text{MLP}(X_{pos})$ (1)
where $\text{Norm}_{L2}$ is the L2 normalization, Concat is the concatenation alongside the channel dimension, and MLP is a multi-layer perceptron with GELU activation [Hendrycks & Gimpel, 2016]. The $X_{rap}$ represents our final relation-aware unified 3D features, which is augmented by the spatial relationship information from Mask3D and the position embeddings.
3.3 ID-Feature Bonding (IFB)
We propose IFB for better referring and grounding in our instruction-following data by improving the connection between object IDs and object features. As shown in Fig. 2 (middle), we first use two identical ID tokens to wrap the object features. Adhering to the causal attention nature of the LLM, this approach links ID information to the object features via the first ID token, and links object information to its ID via the second ID token. Secondly, we propose a post-vision order, which places the vision tokens at the end of the input sequence, closer to the answer tokens generated by the model. This approach mitigates attention deviation from the answer tokens to the ID-Feature pairs, a problem caused by their relative token distance and rotary position embeddings [Su et al., 2024; Ma et al., 2023], while reinforcing the visual information for improved answer generation. The post-vision order is structured as: [$<$System tokens$>$, $<$Instruction tokens$>$, $<$Vision tokens$>$, $<$Answer tokens$>$], where $<$Vision tokens$>$ comprises the ID tokens and object feature tokens.
4 Robust Instruction Generation (RIG)
4.1 Adversarial Data Generation
@@
The Adversarial data is designed to challenge the model’s discriminative capabilities by introducing adversarial or negative samples, ranging from the object-level to the scene-level. It features both category-based identification tasks and expression-based reasoning challenges. As shown in Fig. 3, we ultimately formulate four novel tasks: Hybrid Object Probing Evaluation (Sec. 4.1.1), Hybrid Referring Object Classification (Sec. 4.1.2), Partial Factual 3D Visual Grounding (Sec. 4.1.3), and Faithful 3D Question Answering (Sec. 4.1.4). Details of each type are as follows:
4.1.1 Hybrid Object Probing Evaluation (HOPE) – Fig. 3(upper left)
@@
To construct a scene-level category-based task, we introduce HOPE, which is inspired by the POPE benchmark [Li et al., 2023] in 2D domain. POPE evaluates the tendency of 2D MLLMs to hallucinate by asking yes/no questions about the presence of one specific object at a time. Building on this, HOPE further incorporates such hallucination challenges into the training stage in the 3D domain, aiming to train our model to be more discriminative. Additionally, HOPE presents a hybrid scenario, introducing greater complexity to further advance the decoupling of memorized positive vision and language pairs. Specifically, given a 3D scene, we ask the model to determine the presence of various randomly specified objects. The objects may or may not be present in the scene, and each existing object might have one or more instances. The model is required to answer “No” when the object is not present in the scene, and answer “Yes” with the object ID of each instance of the object when it exists. As shown in Fig. 3 (upper left), the question combines multiple objects, separated by semicolons (“;”), and the answer combines responses for each object, also separated by semicolons. This structure creates a challenging setting that involves hybrid recognition of both positive and negative object presence, combined with multi-instance object localization.
4.1.2 Hybrid Referring Object Classification (HROC) – Fig. 3(upper right)
@@
Referring Object Classification [You et al., 2023] evaluates a model’s ability to understand a referred region in 2D domain, focusing on a classification problem by “Region-in Text-out” format. Our HROC dataset extends this task into the training data for 3D domain to create an object-level category-based task, by incorporating adversarial and hybrid challenges. Specifically, in a 3D scene, we randomly create hybrid positive and negative ID-Category pairs to form our questions, as illustrated in Fig. 3 (upper right). A positive pair consists of a valid object ID and the ground truth category. The bounding box of the corresponding object ID must overlap with one ground truth bounding box, and the category of the overlapping object is defined as the ground truth category. A negative pair includes a valid object ID and a randomly selected category that is present in the scene but not the ground truth category to serve as an adversarial challenge. The model is required to answer “Yes” for positive pairs and “No” with the correct category for negative pairs. The pairs and corresponding answers are separated by semicolons (“;”).
4.1.3 Partial Factual 3D Visual Grounding (PF-3DVG) – Fig. 3(lower left)
@@
Our PF-3DVG introduces a scene-level expression-based task, featuring three types of data in 3DVG: unfactual data, partially factual data, and factual data. For unfactual data, given a 3D scene, we randomly select a reference from Sr3D+ [Achlioptas et al., 2020] where the indicated object does not exist in the scene. The model is required to answer “No” when prompted with the question, as shown in the first example of Fig. 3 (lower left). For partial factual data, given a reference from Sr3D+ and its corresponding 3D scene, we randomly switch the described spatial relationship with a different one based on the predefined template of Sr3D+. For example, as shown in the second example of Fig. 3 (lower left), we change the original reference “the pillow lying on the couch” to “the pillow under the couch”. In this case, it is still possible for human annotators to ground the target based on this partial factual information, but this introduces an adversarial challenge since the information is not completely accurate. Therefore, we require the model to retify the information and answer “It is ‘lying on’” while providing the grounding result (object ID). Notably, we only use references whose target object has no distractors sharing the same category, ensuring that the partial factual information is still informative enough for grounding the target and does not lead to ambiguity. For factual data, we randomly augment the spatial relationship with its synonym to improve diversity. For example, the synonym of “below” can be “under”, “beneath”, or “underneath”.
4.1.4 Faithful 3D Question Answering (3DFQA) – Fig. 3(lower right)
@@
The original 3D Question Answering (QA) task [Azuma et al., 2022] includes only positive samples, which can potentially lead to the model memorizing fixed combinations of 3D scenes and QA pairs. To address this, we propose Faithful 3D Question Answering, a scene-level expression-based task which incorporates both negative and positive examples with an additional grounding requirement. To construct negative samples, we first sample a QA pair and collect the related objects that are mentioned in the question or the target objects of the answer from Azuma et al. [2022]. Then, we randomly select a 3D scene that lacks those related objects. A new instruction is added to the question: “If you can, answer the question… and provide all the IDs…” as illustrated in Fig. 3 (lower right). In this case, the model must faithfully answer “No” based on the absence of related objects in the 3D scene and must not provide any object IDs, demonstrating its reliance on the scene for making decisions. For positive samples, directly taken from Azuma et al. [2022], the model must answer the question while faithfully grounding its “evidence” for the answer, i.e., providing the IDs of the related objects. Therefore, the model trained on our 3DFQA dataset is forced to generalize beyond memorization, learning to respond faithfully to both positive and negative samples.
4.2 Diverse Data Generation

### v2

@@ Abstract
Recent advancements in 3D Large Language Models (3DLLMs) show their potential to build general-purpose agents in the 3D real world, yet challenges remain due to the lack of high-quality robust instruction-following data, leading to limited discriminative power and generalization of 3DLLMs. In this paper, we introduce Robin3D, a powerful 3DLLM trained on large-scale instruction-following data generated by our novel data engine, Robust Instruction Generation (RIG) engine. RIG generates two key instruction data: 1) the Adversarial Instruction-following data, which features mixed negative and positive samples to enhance the model’s discriminative understanding. 2) the Diverse Instruction-following data, which contains various instruction styles to enhance model’s generalization. As a result, we construct 1 million instruction-following data, consisting of 344K Adversarial samples, 508K Diverse samples, and 165K benchmark training set samples. To better handle these complex instructions, Robin3D further integrates an improved vision projector and enhanced sequence organization. Notably, we achieve a 7.8% improvement in the grounding task (Multi3DRefer) and a 6.9% improvement in the captioning task (Scan2Cap).
@@ 1 Introduction
Spatial Intelligence [29] refers to the ability of AI to understand the 3D world and reason within 3D space. Related ideas, such as Embodied AI [13] and Robotic Agent [4], express a similar aim to build general-purpose assistants in the 3D real world. To achieve this goal, researchers have drawn inspiration from the success of 2D Multimodal Large Language Models (MLLMs) [33, 47] and have started exploring the potential of 3D Large Language Models (3DLLMs) [17, 7, 44, 20].
@@
Figure 1: Robin3D surpasses previous SOTA on all 12 benchmarks by training on our RIG-generated 1 million data.
Instruction-following tuning [33, 31, 32] in MLLMs refers to training the LLM to execute natural language commands by integrating both textual and visual information. In contrast to the versatile image-text pairs employed for training 2D MLLMs, collecting 3D instruction-following data for 3DLLM remains a significant challenge. Although existing works have made progress [17, 21, 9] in generating more instruction data, they still lack robustness in two aspects: 1) Most of the existing instruction data consist of positive pairs, lacking adversarial or negative samples. Therefore, a concern remains that models trained on such data tend to be less discriminative, as they risk overfitting to the positive pairs and are more likely to hallucinate positive responses, regardless of the input. 2) Current instruction data lack diversity in language styles, as human annotators or generative models [37, 43] are typically asked to follow fixed instructions when describing objects [6, 10], or the data is generated using predefined templates [1], which may limit models’ generalizability.
To address these challenges, we introduce Robin3D, a robust and powerful 3D Large Language Model tuned on large scale instruction-following data generated by our novel data engine, Robust Instruction Generation (RIG) engine. Specifically, RIG is designed to generate two types of data: Adversarial Instruction-following data, which is characterized by mixing adversarial or negative samples with positive ones. This process decouples the potential memorized positive pairs in the training set, leading to a more discriminative understanding of individual objects and instructions. To present a comprehensive adversarial dataset, we cover both object-level and scene-level instructions, from category-based identification problems to expression-based reasoning challenges, resulting in four new tasks. Diverse Instruction-following data, which first comprehensively collects various types of instructions from existing studies or transforms current tasks into instruction-following format. To harness the powerful in-context learning capability of LLMs, we use ChatGPT [37] to diversify the language styles of the instructions by crafting specific prompt engineering tailored to each task. Combining these with the original training sets of current benchmarks, we construct 1 million instruction-following samples, with approximately 344K adversarial data, 508K diverse data, and 165K benchmark data.
We choose Chat-Scene [19] as the bedrock of Robin3D’s structure due to its efficient extraction of 2D-3D features and its unified approach for object referring and grounding via object IDs. To better handle challenging data from RIG, which contains complex referring and grounding requirements, we incorporate an enhanced vision projector and a more informative way to organize the sequence of tokens. Specifically, a Relation-Augmented Projector is proposed to enrich object-centric features with scene-level context. Additionally, we strengthen the connection between object IDs and object features by wrapping features with identical ID tokens and reinforcing this link via a post-vision order.
In sum, we introduce Robin3D, a powerful 3DLLM trained on robust instruction-following data generated by our novel data engine, RIG. To better handle our generated data, we further involve improvements on the vision projector and sequence organization. As shown in Fig. 1, Robin3D surpasses previous SOTA on all benchmarks without the need for task-specific fine-tuning, different from previous approaches [17, 7].
@@ 2 Related Work
Recent advancements in 3D vision-language (3D-VL) learning [6, 1, 3, 10, 25] have focused on bridging the gap between 3D scene understanding and natural language. Similar to 2D domain [28, 26, 23, 47, 24, 2, 48, 27], tasks like 3D Visual Grounding [6, 49, 1], 3D Dense Captioning [10], and 3D Question Answering [3, 36] have been proposed to evaluate models’ ability to understand human instructions related to 3D objects. Early methods focus on single-task models, such as EDA [45] for grounding and Vote2Cap-DETR++ [8] for captioning. Some studies also develop unified models capable of handling multiple 3D scene-language tasks. Approaches like 3DJCG [5] and D3Net [42] leverage task synergies, while 3D-VisTA [51], 3D-VLP [22] and PQ3D [52] introduce pre-training techniques and unified representations to align 3D vision features with language. However, their dependence on task-specific heads restricts their flexibility for more generalized user-assistant interactions.
@@
Following the success of 2D MLLM [47, 48, 33, 31, 32], researchers begin to explore MLLM in 3D domain (3DLLM). Models like PointLLM [46] and Imagebind-LLM [15] show strong performance in object-level tasks by mapping 3D data into LLMs. However, they face difficulties in handling scene-level reasoning. 3D-LLM [17] incorporates positional embeddings and location tokens, and Oryx [34] offers a solution to support multi-view arbitrary resolution images. However, their reliance on 2D encoders limits the ability to fully capture 3D spatial structures. Models such as LL3DA [7], Chat-3D [44], LEO [21], and Scene-LLM [14] have made progress in improving scene-level dialogue capabilities, showing promising results in question-answering and captioning tasks. However, their insufficient visual grounding capability limits their application in Embodied AI or Robotic Agents, which require precise object localization and manipulation following human instruction. To further enhance grounding abilities, Grounded 3D-LLM [9] introduces referent tokens and the contrastive learning to unify grounding and textual responses. Similarly, Chat-3D v2 [20] proposes the use of object identifiers (object IDs) for referring and grounding. Building on Chat-3D v2[20], Chat-Scene [19] further incorporates DINO v2 [38] to provide strong multi-view, object-centric 2D representations. Despite these advancements, current 3D LLMs, which are trained solely on positive 3D vision-language pairs and template-based instructions, suffer from suboptimal generalization and a potential for overfitting.
3 Approach
@@ 3.1 Preliminary
As a strong baseline for our Robin3D structure, Chat-Scene [19] demonstrates commendable proficiency across multiple benchmarks, attributed to its efficient feature extraction and the use of object IDs for object referring and grounding. They use the pre-trained Mask3D [40] to predict 3D masks for each object. Based on these masks, the point clouds for each object are sampled, normalized, and processed by the pre-trained Uni3D [50] to extract unified, object-centric 3D features. Additionally, 2D masks projected from the 3D masks are used to sample and average 2D features, which are extracted by the pre-trained DINO v2 [39] from multi-view images of each object. Special tokens ${<\texttt{OBJ}i>}_{i=1...n}$ are added to the vocabulary as object IDs and interleaved with the above 3D and 2D object features in the input sequence of LLM to indicate each object for referring in the input and grounding in the output.
However, Chat-Scene only trains benchmark data that are all positive pairs and are labeled or generated by fixed formats, leading to be less discriminative. We detail our generation of robust instruction data in the following and then introduce our improvements on the vision projector and sequence organization to better handle our robust data.
3.2 Robust Instruction Generation (RIG)
3.2.1 Adversarial Data Generation
@@
The Adversarial data is designed to challenge the model’s discriminative capabilities by introducing adversarial or negative samples, ranging from the object-level to the scene-level. It features both category-based identification tasks and expression-based reasoning challenges. As shown in Fig. 2, we ultimately formulate four novel tasks: Hybrid Object Probing Evaluation, Hybrid Referring Object Classification, Partial Factual 3D Visual Grounding, and Faithful 3D Question Answering. Details are as follows:
Hybrid Object Probing Evaluation (HOPE) – Fig. 2(upper left)
@@
To construct a scene-level category-based task, we introduce HOPE, which is inspired by the POPE benchmark [30] in 2D domain. POPE evaluates the tendency of 2D MLLMs to hallucinate by asking yes/no questions about the presence of one specific object at a time. Building on this, HOPE further incorporates such hallucination challenges into the training stage in the 3D domain, aiming to train our model to be more discriminative. Additionally, HOPE presents a hybrid scenario, introducing greater complexity to further advance the decoupling of memorized positive vision and language pairs. Specifically, given a 3D scene, we ask the model to determine the presence of various randomly specified objects. The objects may or may not be present in the scene, and each existing object might have one or more instances. The model is required to answer “No” when the object is not present in the scene, and answer “Yes” with the object ID of each instance of the object when it exists. As shown in Fig. 2 (upper left), the question combines multiple objects, separated by semicolons (“;”), and the answer combines responses for each object, also separated by semicolons. This structure creates a challenging setting that involves hybrid recognition of both positive and negative object presence, combined with multi-instance object localization.
Hybrid Referring Object Classification (HROC) – Fig. 2(upper right)
@@
Referring Object Classification [47] evaluates a model’s ability to understand a referred region in 2D domain, focusing on a classification problem by “Region-in Text-out” format. Our HROC dataset extends this task into the training data for 3D domain to create an object-level category-based task, by incorporating adversarial and hybrid challenges. Specifically, in a 3D scene, we randomly create hybrid positive and negative ID-Category pairs to form our questions, as illustrated in Fig. 2 (upper right). A positive pair consists of a valid object ID and the ground truth category. The bounding box of the corresponding object ID must overlap with one ground truth bounding box, and the category of the overlapping object is defined as the ground truth category. A negative pair includes a valid object ID and a randomly selected category that is present in the scene but not the ground truth category to serve as an adversarial challenge. The model is required to answer “Yes” for positive pairs and “No” with the correct category for negative pairs. The pairs and corresponding answers are separated by semicolons (“;”).
Partial Factual 3D Visual Grounding (PF-3DVG) – Fig. 2(lower left)
@@
Our PF-3DVG introduces a scene-level expression-based task, featuring three types of data in 3DVG: unfactual data, partially factual data, and factual data. For unfactual data, given a 3D scene, we randomly select a reference from Sr3D+ [1] where the indicated object does not exist in the scene. The model is required to answer “No” when prompted with the question, as shown in the first example of Fig. 2 (lower left). For partial factual data, given a reference from Sr3D+ and its corresponding 3D scene, we randomly switch the described spatial relationship with a different one based on the predefined template of Sr3D+. For example, as shown in the second example of Fig. 2 (lower left), we change the original reference “the pillow lying on the couch” to “the pillow under the couch”. In this case, it is still possible for human annotators to ground the target based on this partial factual information, but this introduces an adversarial challenge since the information is not completely accurate. Therefore, we require the model to retify the information and answer “It is ‘lying on’” while providing the grounding result (object ID). Notably, we only use references whose target object has no distractors sharing the same category, ensuring that the partial factual information is still informative enough for grounding the target and does not lead to ambiguity. For factual data, we randomly augment the original spatial relationship with its synonym to improve diversity. For example, the synonym of “below” can be “under”, “beneath”, or “underneath”.
Figure 3: Pipeline to generate our Diverse Instruction data by the in-context learning of ChatGPT.
Figure 4: The one-shot examples for ChatGPT to rephrase the instruction-following data.
Faithful 3D Question Answering (3DFQA) – Fig. 2(lower right)
@@
The original 3D Question Answering (QA) task [3] includes only positive samples, which can potentially lead to the model memorizing fixed combinations of 3D scenes and QA pairs. To address this, we propose Faithful 3D Question Answering, a scene-level expression-based task which incorporates both negative and positive examples with an additional grounding requirement. To construct negative samples, we first sample a QA pair and collect the related objects that are mentioned in the question or the target objects of the answer from [3]. Then, we randomly select a 3D scene that lacks those related objects. A new instruction is added to the question: “If you can, answer the question… and provide all the IDs…” as illustrated in Fig. 2 (lower right). In this case, the model must faithfully answer “No” based on the absence of related objects in the 3D scene and must not provide any object IDs, demonstrating its reliance on the scene for making decisions. For positive samples, directly taken from [3], the model must answer the question while faithfully grounding its “evidence” for the answer, i.e., providing the IDs of the related objects. Therefore, the model trained on our 3DFQA dataset is forced to generalize beyond memorization, learning to respond faithfully to both positive and negative samples.
3.2.2 Diverse Data Generation